A chemist needs 10 L of a 50% alcohol solution. She must mix a 40% solution and a 65% solution. How many liters of each required to obtain what she needs?
Answer by stanbon(57387)

A chemist needs 10 L of a 50% alcohol solution. She must mix a 40% solution and a 65% solution. How many liters of each required to obtain what she needs?
----------------------------------------------------
Equation:
alcohol + alcohol = alcohol
0.40x + 0.65(10-x) = 0.50*10
Multiply thru by 100 to get:
40x + 65(10-x) = 50*10
40x + 650 - 65x = 500
-25x = -150
x = 6 L (amt. of 40% solution in the mixture)
10-x = 4 L (amt. of 65% solution in the mixture)
================================================================
Cheers
Stan H.